the problem says to write the equation in standard form and graph it
x^2+y^2+14y= -13

Found 2 solutions by Theo, richwmiller:
Answer by Theo(13342)

This looks like the equation of a circle.

Standard form of the equation of a circle is:

(x-h)^2 + (y-k)^2 = r^2

Your equation is:

x^2+y^2+14y = -13

you isolate the x terms and y terms together to get:

(x^2) + (y^2 + 14y) = -13

You complete the squares on your x^2 term to get (x-0)^2 = x^2

You complete the squares on your (y^2 + 14y) term to get (y+7)^2 = 6^2 + 14y + 49

You add 49 to the right side of your equation, and your equation becomes:

(x-0)^2 + (y+7)^2 = 36

The center of your circle is (h,k) which equals (0,-7).

To graph this equation you have to solve for y.

Subtract (x-0)^2 from both sides of this equation to get:

(y+7)^2 = -(x-0)^2 + 36

Take the square root of both sides of this equation to get:

y+7 = +/- sqrt(-(x-0)^2 + 36)

Subtract 7 from both sides of this equation to get:

y = +/- (sqrt(-(x-0)^2 + 36)) - 7

A graph of this equation looks like this:



You can see that the center of this circle is at (0,-7) because when x = 0, the top of the circle = (0,-1) and the bottom of the circle = (0,-13). Each one of these points is 6 units away from (0,-7). When y = -7, the left side of the circle = (-6,-7) and the right side of the circle = (6,-7). Each one of these is 6 units away from the center of the circle at (0,-7).

Standard form of the equation is:

(x-0)^2 + (y+7)^2 = 36

Graph of the equation is as shown above, using the formula as shown below:

y = +/- (sqrt(-(x-0)^2 + 36)) - 7 which is really 2 equations as shown below:

y = + (sqrt(-(x-0)^2 + 36)) - 7
y = - (sqrt(-(x-0)^2 + 36)) - 7




Answer by richwmiller(17219)

Just curious.
Have you looked for similar problems before posting?
Have you looked for lessons before posting?